# Kaluguri Yashaswini*, Anshu Arora* and Satish Mulleti

### v1

@@ I INTRODUCTION
Sampling and quantization are fundamental to digital signal processing, enabling continuous-time signals to be represented, stored, and transmitted efficiently. Quantization is typically uniform (UQ) or non-uniform (NUQ). While UQ is simple, it is suboptimal when the signal amplitudes are not uniformly distributed, whereas NUQ adapts resolution to the statistics of the data and yields improved rate–distortion performance [1, 2, 3, 4, 5].
As in the quantization, unlike uniform sampling [6], event-driven strategies such as non-uniform sampling (NUS) [7, 8, 9] and time encoding machines (TEMs) [10, 11, 12, 13, 14] offer more efficient representations by adapting sampling locations to the signal variations. In particular, the integrate-and-fire TEM (IF–TEM) stands out for its simplicity and low power consumption [15, 16, 17, 18, 19]. By integrating the input and recording firing times whenever a threshold is reached, the IF–TEM produces time encodings whose inter-event intervals adapt naturally to the input dynamics.
Prior work [15] analyzed UQ of IF–TEM intervals and showed that the induced quantization error has the same upper bound as quantizing amplitudes in NUS. However, since the firing intervals are rarely uniformly distributed, UQ is inefficient for IF–TEM outputs. This motivates the design of NUQ schemes tailored to the statistics of time encodings.
@@ II PROBLEM FORMULATION
We consider the input signal as a real-valued, stationary, ergodic, and bandlimited stochastic process $\{f(t)\}_{t\in\mathbb{R}}\in\mathcal{B}_{\Omega_{0},c}$, with spectral support in $\Omega_{0},\Omega_{0}]$ and bounded amplitude, that is, $|f(t)|\leq c$. For any $f(t)\in\mathcal{B}_{\Omega_{0},c}$, its Fourier transform is defined as $F(\Omega)=\int_{-\infty}^{\infty}f(t)e^{\mathrm{j}\Omega t}\,\mathrm{d}t,$ and satisfies $F(\Omega)=0$ for all $|\Omega|>\Omega_{0}$. In addition, the amplitude of these signals is bounded such that $|f(t)|\leq c$. It is well-known by the Shannon-Nyquist sampling theory [20],[6] that any $f(t)\in\mathcal{B}_{\Omega_{0},c}$ can be exactly reconstructed from its uniform samples $\{f(nT_{s})\}$, provided that the sampling period obeys $T_{s}<T_{Nyq}=\frac{\pi}{\Omega_{0}}.$ The discrete representation of the samples $f(nT_{s})$ is then quantized for further processing. Alternatively, the signal can also be reconstructed from the NUS set $\{f(t_{n}),t_{n}\}_{n\in\mathbb{Z}}$ provided that $t_{n+1}-t_{n}<T_{Nyq}$ [21]. However, unlike the uniform sampling, here, both amplitudes and sampling locations have to be quantized.
@@
Once the intervals $\{T_{n}\}$ are given, the signal $f(t)$ can be perfectly reconstructed by applying an iterative algorithm, provided that $T_{n}<T_{\text{Nyq}},\ \forall n$ [15]. We must select $\kappa$ and $\Delta$ according to (3) such that this condition is satisfied. Without loss of generality, we set $\kappa=1$ and keep $\Delta$ as a free variable.
@@ III-A Theoretical Results
Since $f(t)$ is bandlimited, it is continuous, and by the Mean Value Theorem for integrals [22, Chapter 6], there exists a point $\tau_{n}\in[t_{n},t_{n+1}]$ such that
@@
By stationarity, the distribution of $f(\tau_{n})$ coincides with the ensemble distribution of $f(t)$ at any fixed time [23, Chapter 9]. Thus,
@@
The probability density function (PDF) of $\mathbf{T}$ induced by the mapping $g(f)=\Delta/(f+b)$ is obtained via the change-of-variables theorem [23, Chapter 5]. Since $g$ is monotone on [c,c][-c,c], we have
@@
Since $\{f(\tau_{n})\}_{n\in\mathbb{N}}$ is stationary and ergodic, and the mapping $g(x)=\tfrac{\Delta}{b+x}$ is measurable and bounded (as ensured by (13)), it follows that the sequence $\{T_{n}\}$ is also stationary and ergodic [24, Chapter 6].
Thus, by Birkhoff’s Ergodic Theorem [24, Chapter 1], for any integrable function $\varphi$,
@@ III-B Practical NUQ Design for IF–TEM
Fig. 3(a) shows the averaged histogram of the signal amplitudes computed over the $100$ independent realizations(signals). This averaging procedure provides an empirical estimate of the ensemble law $P_{\mathbf{F}}(f)$ [23, chapter 7] of the bandlimited stochastic process. Fig. 3(b) presents an overlay of the theoretically derived PDF of the interspike intervals $P_{\mathbf{T}}(T)$, obtained from (4), and the observed distribution of $\{T_{n}\}$ estimated from the simulated signals as for the amplitudes. The close agreement between the theoretical and simulated distributions validates the analytical derivation of the interspike interval ensemble distribution and confirms the accuracy of the IF–TEM simulations.
Motivated by this characterization, we now design non-uniform quantizers for the firing intervals. In principle, NUQ can be implemented directly using optimization-based methods such as the Lloyd–Max algorithm, or indirectly by applying a suitable transformation followed by uniform quantization. The latter approach is particularly well-suited for the present application, since power-law based companding schemes are inherently more robust to distribution mismatches and outliers than direct Lloyd–Max designs. Accordingly, we employ a companding-based strategy similar to [5, 25], in which the distribution is first nonlinearly transformed as

### v2

@@ I INTRODUCTION
Sampling and quantization are fundamental to digital signal processing, enabling continuous-time signals to be represented, stored, and transmitted efficiently. Quantization is typically uniform (UQ) or non-uniform (NUQ). While UQ is simple, it is suboptimal when the signal amplitudes are not uniformly distributed, whereas NUQ adapts resolution to the statistics of the data and yields improved rate–distortion performance [gersho2012vector, companding, max_lloyd, hila2022time, escort1].
As in the quantization, unlike uniform sampling [jerri1977shannon], event-driven strategies such as non-uniform sampling (NUS) [marvastinonuniform, irregsampling, aldroubi2001nonuniform] and time encoding machines (TEMs) [lazar2005time, gontier2014sampling, VBIFTEM, omar2024adaptive, hilaTEM] offer more efficient representations by adapting sampling locations to the signal variations. In particular, the integrate-and-fire TEM (IF–TEM) stands out for its simplicity and low power consumption [PerfectRecovery, compressediftem, lowpower, mulleti2023powerefficientsampling, hilton2021guaranteed]. By integrating the input and recording firing times whenever a threshold is reached, the IF–TEM produces time encodings whose inter-event intervals adapt naturally to the input dynamics.
Prior work [PerfectRecovery] analyzed UQ of IF–TEM intervals and showed that the induced quantization error has the same upper bound as quantizing amplitudes in NUS. However, since the firing intervals are rarely uniformly distributed, UQ is inefficient for IF–TEM outputs. This motivates the design of NUQ schemes tailored to the statistics of time encodings.
@@ II PROBLEM FORMULATION
We consider the input signal as a real-valued, stationary, ergodic, and bandlimited stochastic process $\{f(t)\}_{t\in\mathbb{R}}\in\mathcal{B}_{\Omega_{0},c}$, with spectral support in $\Omega_{0},\Omega_{0}]$ and bounded amplitude, that is, $|f(t)|\leq c$. For any $f(t)\in\mathcal{B}_{\Omega_{0},c}$, its Fourier transform is defined as $F(\Omega)=\int_{-\infty}^{\infty}f(t)e^{\mathrm{j}\Omega t}\,\mathrm{d}t,$ and satisfies $F(\Omega)=0$ for all $|\Omega|>\Omega_{0}$. In addition, the amplitude of these signals is bounded such that $|f(t)|\leq c$. It is well-known by the Shannon-Nyquist sampling theory [nyquist1928certain],[jerri1977shannon] that any $f(t)\in\mathcal{B}_{\Omega_{0},c}$ can be exactly reconstructed from its uniform samples $\{f(nT_{s})\}$, provided that the sampling period obeys $T_{s}<T_{Nyq}=\frac{\pi}{\Omega_{0}}.$ The discrete representation of the samples $f(nT_{s})$ is then quantized for further processing. Alternatively, the signal can also be reconstructed from the NUS set $\{f(t_{n}),t_{n}\}_{n\in\mathbb{Z}}$ provided that $t_{n+1}-t_{n}<T_{Nyq}$ [feichtinger1994theory]. However, unlike the uniform sampling, here, both amplitudes and sampling locations have to be quantized.
@@
Once the intervals $\{T_{n}\}$ are given, the signal $f(t)$ can be perfectly reconstructed by applying an iterative algorithm, provided that $T_{n}<T_{\text{Nyq}},\ \forall n$ [PerfectRecovery]. We must select $\kappa$ and $\Delta$ according to (3) such that this condition is satisfied. Without loss of generality, we set $\kappa=1$ and keep $\Delta$ as a free variable.
@@ III-A Theoretical Results
Since $f(t)$ is bandlimited, it is continuous, and by the Mean Value Theorem for integrals [rudin1976principles, Chapter 6], there exists a point $\tau_{n}\in[t_{n},t_{n+1}]$ such that
@@
By stationarity, the distribution of $f(\tau_{n})$ coincides with the ensemble distribution of $f(t)$ at any fixed time [papoulis2002probability, Chapter 9]. Thus,
@@
The probability density function (PDF) of $\mathbf{T}$ induced by the mapping $g(f)=\Delta/(f+b)$ is obtained via the change-of-variables theorem [papoulis2002probability, Chapter 5]. Since $g$ is monotone on [c,c][-c,c], we have
@@
Since $\{f(\tau_{n})\}_{n\in\mathbb{N}}$ is stationary and ergodic, and the mapping $g(x)=\tfrac{\Delta}{b+x}$ is measurable and bounded (as ensured by (13)), it follows that the sequence $\{T_{n}\}$ is also stationary and ergodic [walters1982introduction, Chapter 6].
Thus, by Birkhoff’s Ergodic Theorem [walters1982introduction, Chapter 1], for any integrable function $\varphi$,
@@ III-B Practical NUQ Design for IF–TEM
Fig. 3(a) shows the averaged histogram of the signal amplitudes computed over the $100$ independent realizations(signals). This averaging procedure provides an empirical estimate of the ensemble law $P_{\mathbf{F}}(f)$ [papoulis2002probability, chapter 7] of the bandlimited stochastic process. Fig. 3(b) presents an overlay of the theoretically derived PDF of the interspike intervals $P_{\mathbf{T}}(T)$, obtained from (4), and the observed distribution of $\{T_{n}\}$ estimated from the simulated signals as for the amplitudes. The close agreement between the theoretical and simulated distributions validates the analytical derivation of the interspike interval ensemble distribution and confirms the accuracy of the IF–TEM simulations.
Motivated by this characterization, we now design non-uniform quantizers for the firing intervals. In principle, NUQ can be implemented directly using optimization-based methods such as the Lloyd–Max algorithm, or indirectly by applying a suitable transformation followed by uniform quantization. The latter approach is particularly well-suited for the present application, since power-law based companding schemes are inherently more robust to distribution mismatches and outliers than direct Lloyd–Max designs. Accordingly, we employ a companding-based strategy similar to [escort1, escort2], in which the distribution is first nonlinearly transformed as
@@ V CONCLUSION
VI APPENDIX
VI-A Comparison to Uniform Quantization
In this section, we present additional simulation results obtained using practical signals. We also compare the performance of uniform sampling (US), using sinc interpolation, with uniform and non-uniform quantization. The US is performed and reconstructed as follows:
$f(t)=\sum_{n=-\infty}^{\infty}f(nT_{s})\ \text{sinc}\left(\frac{t-nT_{s}}{T_{s}}\right)$ (17)
where $\text{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$, and $T_{s}\leq\frac{\pi}{\Omega_{0}}$. In Fig. 6, we compare the performance of US with amplitude UQ and NUQ, for the same class of signals, as well as the same number of samples, as presented in Section IV. We observe that TEM-NUQ performs significantly better than both these schemes, and we observe a negligible difference between UQ and NUQ for the US scheme.
VI-B Practical Application
In this section, we consider signals from practical applications. Specifically, we analyze an experimentally acquired ultrasonic guided wave signal reported in [Patil2024], obtained from a steel pipe instrumented with circumferential piezoelectric thickness shear mode transducers. Torsional guided wave responses recorded at the receivers are compared with a baseline signal from a healthy pipe to identify defects through signal deviations. The received signal is used in this work to evaluate the performance of the proposed TEM-NUQ scheme for the ultrasonic guided wave signals.
The following results in Fig. 7 depict the reconstruction error with the number of bits, for this signal, compared with the US scheme as well. We observe a better reconstruction error across all three sampling strategies in the case of NUQ, as designed using the transformation presented in Theorem 1. We would like to point out that we are using the same number of samples for all three strategies. The parameters used are $b=1.67,\Delta=3.03\times 10^{-6}$ for the TEM strategy, and the sampling frequency $F_{s}=6\times 10^{5}$ Hz.
Figure 7: NMSE comparison of TEM, NUS, US under uniform and non-uniform quantization.
(a) Non-Uniform quantization in TEM with 4 bits ($\text{NMSE}=-29.952\,\mathrm{dB}$).
(b) Uniform quantization in TEM with 4 bits ($\text{NMSE}=-22.789\,\mathrm{dB}$).
Figure 8: Reconstruction error comparison between TEM-NUQ and TEM-UQ for the ultrasonic guided waves.
As depicted in Fig. 8, at a quantization resolution of $4$ bits, TEM-NUQ achieves the lowest reconstruction error among all considered schemes, providing an improvement of approximately $7$ dB compared to TEM-UQ. Here, TEM-NUQ employs non-uniformly quantized time encodings matched to the signal statistics, whereas TEM-UQ uses uniformly quantized time encodings. These results validate the effectiveness of the proposed NUQ design on practical ultrasonic guided wave signals.